### v1

@@ 3.1 Asymptotic minors
Without loss of generality assume $i<j$. Let $S$ denote the $(a_{i},a_{j})$-subpath of $P$. Observe that $|A_{r}\left(S\right)|\geq 10K+9$. Let $X_{i}$ and $X_{j}$ denote the $(a_{i},b_{i})$-subpath of $P$. Due to 5.2, we have that $|A_{r}\left(X_{i}\right)|+|A_{r}\left(X_{j}\right)|\leq 6$. Now consider the case when both $b_{i}$ and $b_{j}$ lies on the path $S$. In this case, $S$ is obtained by concatenating $X_{i}$, $R_{ij}$ and $X_{j}$. Hence, we have the following.
$|A_{r}\left(X_{i}\right)|+|A_{r}\left(R_{ij}\right)|+|A_{r}\left(X_{j}\right)|$ $\geq|A_{r}\left(S\right)|$
$|A_{r}\left(R_{ij}\right)|+6$ $\geq 10K+9$
$|A_{r}\left(R_{ij}\right)|$ $\geq 10K+3$
In the remaining cases i.e. when $b_{i}$ or $b_{j}$ does not lie on the path $S$, similar arguments prove the claim. ∎
The above claim implies that the $(b_{i},b_{j})$-subpath of $P$, with $i\neq j$, necessarily contains many anti-chain vertices.
Claim 5.4.
For an integer $i\in[t]$, let $c\in V(P)$ be a vertex such that the $(c,b_{i})$-subpath of $P$ contain at least $32K+$ 3 antichain vertices. Then for any vertex $d\in V(Q_{i})$, $\mathsf{d}\left(c,d\right)\geq K$.
@@
Suppose $\mathsf{d}\left(c,d\right)<K$. Then $\mathsf{d}\left(c,b_{i}\right)\leq 2k+1+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,c\right)|$. But this contradicts Proposition 2 which implies that $\mathsf{d}\left(c,b_{i}\right)\geq 2K+2+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,c\right)|$. ∎
For each $i\in[t]$, let $Q^{\prime}_{i}$ denote the maximal $(b_{i},c_{i})$-subpath of $Q_{i}$ such that $\mathsf{d}\left(Q^{\prime}_{i},Q_{j}\right)\geq K$ for some $i\neq j,j\in[t]$. Now we prove a lower bound on the length of $Q^{\prime}_{i}$, for each $i\in[t]$. We prove the following final claim before providing the explicit constructions of the branch paths and branch sets.
Claim 5.5.
For each $i\in[t]$, the path $Q^{\prime}_{i}$ has length at least $18K+$ 1.
@@
Assume for contradiction that there exists an integer $i\in[t]$ such that the length of $Q^{\prime}_{i}$ is at most $8K$. Let $c^{\prime}_{i}$ be the vertex adjacent to $c_{i}$ in $Q_{i}$ which is not a vertex of $Q^{\prime}_{i}$. Let $j\neq i$ be the minimum index such that $c^{\prime}_{i}$ lies within a distance of at most $2K$ from a vertex $d_{j}\in V(Q_{j})$. Let $T_{ij}$ denote a $(c^{\prime}_{i},d_{j})$-isometric path in $G$. Now, observe that the path $X=Q^{\prime}_{i}\oplus\{c^{\prime}_{i}\}\oplus T_{ij}\oplus Q^{\prime}_{j}$ is a $(b_{i},b_{j})$-path. Since $\mathsf{d}\left(b_{i},b_{j}\right)$ is at most the length of $X$ we have that
$\mathsf{d}\left(b_{i},b_{j}\right)\leq 10K+|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,b_{i}\right)|$
Let $R_{ij}$ denotes the $(b_{i},b_{j})$-subpath of $P$. From 5.3 it follows that $|A_{r}\left(R_{ij}\right)|\geq 10K+3$. Now from Proposition 2 it follows that
$\mathsf{d}\left(b_{i},b_{j}\right)\geq 10K+2+|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,b_{i}\right)|$
which is a contradiction. ∎
@@
Constructing branch paths: Let $\mathcal{Q}=\{Q^{\prime}_{i}\colon i\in[t]\}$. Let $X$ be the $(b_{i},b_{i+1})$-subpath of $P$ and let $D_{i}$ denote the maximum subpath of $X$ such that both components of the graph $E(X)\setminus E(D_{i})$ contain $24K+$ 2 antichain vertices. Observe that $D_{i}$ always exists (due to 5.3). Let $\mathcal{D}=\{D_{i}\colon i\in[t-1]\}\}$. Moreover, each $D_{i}$ contains at least $K+3$ antichain vertices. We shall use this fact later.
Claim 5.6.
For each $i\in[t-1]$, the length of $D_{i}$ is at least $K+3$.
We now show that the paths in $\mathcal{Q}\cup\mathcal{D}$ can be used as branch paths.
Claim 5.7.
@@
If $X,Y\in\mathcal{C}$, then we are done by the definition of $\mathcal{Q}$. Suppose $X$ is $Q^{\prime}_{i}$ for some $i\in[t]$ and $Y=D_{j}$ for some $j\in[t-1]$. Assume that $\mathsf{d}\left(X,Y\right)\leq K$ and $y\in Y$ be a vertex such that $\mathsf{d}\left(y,X\right)<K$. Note that $y\in V(P)$ and the $(b_{i},y)$-subpath of $P$, say $P^{\prime}$, contains at least $32K+$ 3 antichain vertices. Hence, we have a contradiction due to 5.4. ∎
Constructing branch sets: For each $i\in[t]$, let $c^{\prime}_{i}$ denote the end-vertex of $Q^{\prime}_{i}$ that is closest to $r$, and let $T_{i}$ denote the $(r,c^{\prime}_{i})$-subpath of $Q_{i}$. Let $T$ denote the set $\bigcup\limits_{i\in I}V(T_{i})$. Clearly, the subgraph induced by $T$ is connected and $T$ will be the branch set corresponding to the universal vertex of $U_{t}$. Below, we construct the branch sets corresponding to the other vertices of $U_{t}$. We shall use the paths in $\mathcal{D}$ (defined above) to construct these branch sets.
For each $i\in[t]$, let $X_{i}$ denote the $(a_{i},b_{i})$-subpath of $P$ and $P_{i}=V(X_{i})\cup\{b^{\prime}_{i}\}$. Observe that the subgraph induced by $P_{i}$ is connected. For $i\in[t-1]$, let $R_{i}$ denote the $(b_{i},b_{i+1})$-subpath of $P$. Let $Y_{1}$ be the vertices of the connected component of $R_{1}\setminus D_{1}$ that contains $b_{1}$. For $i\in[2,t-1]\}$, let $Y_{i}$ be the union of the vertices of the connected components of $E(R_{i-1})\setminus E(D_{i-1})$ and $E(R_{i+1})\setminus E(D_{i+1})$ that contain $b_{i}$. Finally, let $Y_{t}$ be the vertices of the connected component of $R_{t-1}\setminus D_{t-1}$ that contains $b_{t}$. For $i\in[t]$, let $Z_{i}=Y_{i}\cup P_{i}$. Note that the subgraph induced by $Z_{i}$ is connected and $Z_{i}\subseteq V(Q^{\prime}_{i})\cup V(P)$. We have the following claims.
Claim 5.8.
@@
Suppose that there exists an integer $i\in[t]$, such that $\mathsf{d}\left(x,y\right)<K$ with $x\in T$ and $y\in Z_{i}$. The definition of $T$ implies that $x\in Q_{j}$ for some $j\in[t]$. 5.5 immediately implies that $y\in V(P)\cap Z_{j}$. Now if $i\neq j$, the $(y,b_{j})$-subpath of $P$ contains at least $34K+$ 3 antichain vertices, which is a contradiction due to 5.4. Now consider the case when $i=j$. Let $c^{\prime}_{i}$ denote the end-vertex of $Q^{\prime}_{i}$ that is different from $b_{i}$. From the definition of $T$ (and $T_{i}$), it follows that $\mathsf{d}\left(r,x\right)\leq\mathsf{d}\left(r,c^{\prime}_{i}\right)$.
Let $d^{\prime}_{i}$ denote the vertex of $Q_{i}$ which is adjacent to $c^{\prime}_{i}$ but does not lie in $Q^{\prime}_{i}$. The definition of $Q^{\prime}_{i}$ implies that there exists an integer $k\in[t]\setminus\{i\}$, such that $\mathsf{d}\left(c^{\prime}_{i},Q^{\prime}_{k}\right)<K$. Let $d^{\prime}_{k}$ denote a vertex of $Q^{\prime}_{k}$ with $\mathsf{d}\left(d^{\prime}_{i},d^{\prime}_{k}\right)<K$. Observe the following:
@@
$\leq 1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,d^{\prime}_{k}\right)|+K+|\mathsf{d}\left(r,d^{\prime}_{i}\right)-\mathsf{d}\left(r,x\right)|+K$
$\leq 4K+1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,y\right)|$

### v2

@@ 3.1 Asymptotic minors
See Figure˜5(b). Without loss of generality, assume that $i<j$. Let $S$ denote the $(a_{i},a_{j})$-subpath of $P$. Observe that by the definition of $a_{i}$ and $a_{j}$, $|A\cap V(S)|\geq{12K}+{5}$. Let $X_{i}$ and $X_{j}$ denote the $(a_{i},b_{i})$-subpath and $(a_{j},b_{j})$-subpath of $P$, respectively. Due to ˜3.1.B, we have that $|A\cap V(X_{i})|\leq 3$, $|A\cap V(X_{j})|\leq 3$. According to Procedure 1, $b_{j}$ always lies in $S$. Hence we have two cases depending on whether $b_{i}$ lies in $S$ or not. First, consider the case when $b_{i}\in V(S)$. In this case, $b_{i}=a_{i}$. Indeed, according to Procedure 1, $b_{i}$ lies on the $(u,a_{i})$-subpath of $P$ and therefore $b_{i}$ lies in $S$ only if $b_{i}=a_{i}$. Therefore, $S$ can be constructed by concatenating $R_{ij}$ and $X_{j}$. Hence, we have the following.
$|A\cap V(R_{ij})|+|A\cap V(X_{j})|$ $\geq|A\cap V(S)|$
$|A\cap V(R_{ij})|+3$ $\geq{12K}+{5}$
$|A\cap V(R_{ij})|$ $\geq{12K}+2$
Consider the case $b_{i}\not\in V(S)$. Observe that $X_{j}$ is a subpath of $S$, and that the subpath $S^{\prime}$ of $S$ obtained by deleting the vertices of $X_{j}$ is a subpath of $R_{ij}$. Hence, $|A\cap V(R_{ij})|\geq|A\cap V(S^{\prime})|\geq|A\cap V(S)|-|A\cap V(X_{j})|\geq 12K+5$.
The above claim implies that the $(b_{i},b_{j})$-subpath of $P$, with $i\neq j$, necessarily contains an antichain of large size. For each $i\in[t]$, let $Q^{\prime}_{i}$ denote the maximal subpath of $Q_{i}$ such that (i) $Q^{\prime}_{i}$ has $b_{i}$ as an end-vertex and (ii) $\mathsf{d}\left(Q^{\prime}_{i},Q_{j}\right)\geq K$ for all $i\neq j,j\in[t]$. Note that $Q^{\prime}_{i}$ is well-defined, as for all $i,j\in[t]$, $\mathsf{d}\left(b_{i},Q_{j}\right)\geq 6K$. (Otherwise, it can be shown, using similar arguments as the ones in ˜3.1.D, that $\mathsf{d}\left(b_{i},b_{j}\right)\leq 12K$, contradicting ˜3.1.C and Proposition˜13. We omit these details as ˜3.1.D already implies that $Q^{\prime}_{i}$ is well-defined for each $i\in[t]$.) Let $c_{i}$ be the end-vertex of $Q^{\prime}_{i}$ which is distinct from $b_{i}$. Now we prove a lower bound on the length of $Q^{\prime}_{i}$, for each $i\in[t]$. We prove the following final claim before providing the explicit constructions of the branch paths and branch sets.
Claim 3.1.D.
For each $i\in[t]$, the path $Q^{\prime}_{i}$ has length at least $5K+1$.
@@
See Figure˜5(c). Assume for contradiction that there exists an integer $i\in[t]$ such that the length of $Q^{\prime}_{i}$ is at most $5K$. Let $c^{\prime}_{i}$ be the vertex adjacent to $c_{i}$ in $Q_{i}$ which is not a vertex of $Q^{\prime}_{i}$. Let $j\neq i$ be the minimum index such that $c^{\prime}_{i}$ has distance strictly less than $K$ from some vertex $d_{j}\in V(Q_{j})$. Let $T_{ij}$ denote a $(c^{\prime}_{i},d_{j})$-isometric path in $G$. Let $d^{\prime}_{j}$ be the vertex of $Q_{j}$ such that $\mathsf{d}\left(r,d^{\prime}_{j}\right)=\max\{\mathsf{d}\left(r,c^{\prime}_{i}\right),\mathsf{d}\left(r,d_{j}\right)\}$. In other words, $d^{\prime}_{j}$ is the vertex of $Q_{j}$ which is either $d_{j}$ or lies at the same distance from $r$ as $c^{\prime}_{i}$. Let $f_{j}$ be the vertex of $Q_{j}$ such that $\mathsf{d}\left(r,f_{j}\right)=\max\{\mathsf{d}\left(r,b_{i}\right),\mathsf{d}\left(r,b_{j}\right)\}$. In other words, $f_{j}$ is the vertex of $Q_{j}$ which is either $b_{j}$ or lies at the same distance from $r$ as $b_{i}$. Observe that
$\mathsf{d}\left(b_{i},b_{j}\right)\leq\mathsf{d}\left(b_{i},c^{\prime}_{i}\right)+\mathsf{d}\left(c^{\prime}_{i},d_{j}\right)+\mathsf{d}\left(d_{j},d^{\prime}_{j}\right)+\mathsf{d}\left(d^{\prime}_{j},f_{j}\right)+\mathsf{d}\left(f_{j},b_{j}\right).$
Recall that $\mathsf{d}\left(b_{i},c^{\prime}_{i}\right)\leq 5K+1$ and $\mathsf{d}\left(c^{\prime}_{i},d_{j}\right)\leq K-1$. By the definition of $d^{\prime}_{j}$ and $f_{j}$, these imply that $\mathsf{d}\left(d^{\prime}_{j},f_{j}\right)\leq 5K+1$ and $\mathsf{d}\left(d_{j},d^{\prime}_{j}\right)\leq K-1$. Finally, the distance between $f_{j}$ and $b_{j}$ is $|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,f_{j}\right)|$. Now we have that
$\mathsf{d}\left(b_{i},b_{j}\right)\leq{12K}+|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,b_{i}\right)|.$
Let $R_{ij}$ denote the $(b_{i},b_{j})$-subpath of $P$. From ˜3.1.C it follows that $|A\cap V(R_{ij})|\geq{12K}+2$. Now from Proposition˜13, it follows that
$\mathsf{d}\left(b_{i},b_{j}\right)\geq{12K}+1+|\mathsf{d}\left(r,b_{j}\right)-\mathsf{d}\left(r,b_{i}\right)|$
which is a contradiction. ∎
The following claim implies that if the subpath between some vertex $c\in V(P)$ and $b_{i}$, for some $i\in[t]$, contain many antichain vertices, then, $c$ must be far away from $Q^{\prime}_{i}$.
Claim 3.1.E.
For an integer $i\in[t]$, let $c\in V(P)$ be a vertex such that the $(c,b_{i})$-subpath of $P$ contains at least $2K+1$ antichain vertices of $A$. Then, for any vertex $d\in V(Q_{i})$, $\mathsf{d}\left(c,d\right)\geq K$.
@@
See Figure˜5(d). Suppose by contradiction that $\mathsf{d}\left(c,d\right)<K$ for some $d\in V(Q_{i})$. Due to Proposition˜13, $\mathsf{d}\left(c,b_{i}\right)\geq 2K+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,c\right)|$. Observe that $\mathsf{d}\left(d,b_{i}\right)\leq K$ would imply $\mathsf{d}\left(c,b_{i}\right)<2K$, contradicting the above. Hence, there must exist a vertex $z\in V(Q_{i})$, such that $\mathsf{d}\left(r,z\right)=\mathsf{d}\left(r,c\right)$ and $\mathsf{d}\left(d,z\right)<K$. Hence, $\mathsf{d}\left(b^{\prime}_{i},c\right)\leq\mathsf{d}\left(b^{\prime}_{i},z\right)+\mathsf{d}\left(z,d\right)+\mathsf{d}\left(c,d\right)<2K+\mathsf{d}\left(b^{\prime}_{i},z\right)\leq 2K+\mathsf{d}\left(r,b^{\prime}_{i}\right)-\mathsf{d}\left(r,z\right)$. Since $b^{\prime}_{i}$ and $b_{i}$ are adjacent, $|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,b^{\prime}_{i}\right)|\leq 1$. Hence, $\mathsf{d}\left(b_{i},c\right)\leq\mathsf{d}\left(b^{\prime}_{i},c\right)+1<2K+1+\mathsf{d}\left(r,b^{\prime}_{i}\right)-\mathsf{d}\left(r,z\right)=2K+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,c\right)|$, a contradiction. ∎
[…]
(a) The shaded triangular shapes indicate the branch sets $Z_{1},Z_{2},\ldots,Z_{t}$. The paths $Q^{\prime}_{1},Q^{\prime}_{2},\ldots,Q^{\prime}_{t}$ and some and subpaths of $P$ (not indicated in this figure) are used as branch paths to construct a $K$-fat minor model of $U_{t}$.
[…]
(b) Construction of $Z_{i}$. Vertices connected by the thick gray lines indicate $Z_{i}$. The dotted lines indicate $Q_{i}$.
Figure 6: Construction of branch sets and branch paths.
We are now ready to construct the $K$-fat minor model of $U_{t}$, in order to provide a contradiction.
@@
Constructing the branch paths and branch sets. Figure˜6 illustrates (on a high level) the construction of branch paths and branch sets of the $K$-fat minor model of $U_{t}$. Let $\mathcal{Q}=\{Q^{\prime}_{i}\colon i\in[t]\}$. Let $R_{i}$ be the $(b_{i},b_{i+1})$-subpath of $P$ and let $D_{i}$ denote the maximum subpath of $R_{i}$ such that both components of the graph $E(R_{i})\setminus E(D_{i})$ contain $5K+2$ antichain vertices of $A$. Observe that $D_{i}$ always exists (due to ˜3.1.C). Let $\mathcal{D}=\{D_{i}\colon i\in[t-1]\}\}$.
Recall that, for $i\in[t]$, $c_{i}$ denotes the end-vertex of $Q^{\prime}_{i}$ that is closest to $r$, and let $T_{i}$ denote the $(r,c_{i})$-subpath of $Q_{i}$. Let $T$ denote the set $\bigcup\limits_{i\in[t]}V(T_{i})$.
Clearly, the subgraph induced by $T$ is connected and $T$ will be the branch set corresponding to the universal vertex of $U_{t}$. Below, we construct the branch sets corresponding to the other vertices of $U_{t}$. We shall use the paths in $\mathcal{D}$ (defined above) to construct these branch sets.
For each $i\in[t]$, let $X_{i}$ denote the $(a_{i},b_{i})$-subpath of $P$ and $P_{i}=V(X_{i})\cup\{b^{\prime}_{i}\}$. Observe that the subgraph induced by $P_{i}$ is connected.
For $i\in[t-1]$, let $R_{i}$ denote the $(b_{i},b_{i+1})$-subpath of $P$. Let $Y_{1}$ be the vertices of the connected component of $R_{1}\setminus D_{1}$ that contains $b_{1}$. For $i\in[2,t-1]$, let $Y_{i}$ be the union of the vertices of the connected components of $R_{i-1}\setminus D_{i-1}$ and $R_{i+1}\setminus D_{i+1}$ that contain $b_{i}$. Finally, let $Y_{t}$ be the vertices of the connected component of $R_{t-1}\setminus D_{t-1}$ that contains $b_{t}$. For $i\in[t]$, let $Z_{i}=Y_{i}\cup P_{i}$. Note that the subgraph induced by $Z_{i}$ is connected and $Z_{i}\subseteq V(Q^{\prime}_{i})\cup V(P)$. To complete the proof, we need to show the following:
  1. 1.
    for each $i\in[t]$, $T\cap Q^{\prime}_{i}\neq\emptyset$;
  2. 2.
    $D_{1}\cap Z_{1}\neq\emptyset,D_{t-1}\cap Z_{t}\neq\emptyset$; and
  3. 3.
  4. 4.
  5. 5.
    for each $i,j\in[t]$ with $i\neq j$, $\mathsf{d}\left(Z_{i},Q^{\prime}_{j}\right)\geq K$,
  6. 6.
    for any $i\in[t]$, $\mathsf{d}\left(T,Z_{i}\right)\geq K$.
  7. 7.
    for any $i,j\in[t]$ with $i<j$, $\mathsf{d}\left(Z_{i},Z_{j}\right)\geq K$,
  8. 8.
    for any $i\in[t-1]$, $\mathsf{d}\left(T,D_{i}\right)\geq K$, and
  9. 9.
    for any $i\in[t-1],j\in[t]$ with $|j-i|\notin\{0,1\}$, we have $\mathsf{d}\left(D_{i},Z_{j}\right)\geq K$.
Items˜1, 2 and 3 follow from definitions. The next claim proves that the branch paths are far apart from each other (Item˜4).
Claim 3.1.F.
@@
If $X,Y\in\mathcal{Q}$, then we are done by the definition of $\mathcal{Q}$. Suppose $X$ is $Q^{\prime}_{i}$ for some $i\in[t]$ and $Y=D_{j}$ for some $j\in[t-1]$. Since any vertex $y\in V(Y)$ is a vertex of $P$, the $(b_{i},y)$-subpath of $P$, say $P^{\prime}$, contains at least $5K+2$ antichain vertices of $A$. Now, ˜3.1.E implies the current claim.
Finally, assume that $X=D_{i}$ for some $i\in[t-1]$. By definition, both components found after deleting $X$ from the $(b_{i},b_{i+1})$-subpath have $5K+2$ antichain vertices of $A$, i.e., have length at least $5K+3$. So, if both $X,Y$ are in $\mathcal{D}$, they are part of $P$, and thus their distance is at least $5K+3\geq K$. Hence, also in this case the claim holds. ∎
The following claim proves Item˜5 (in a slightly stronger sense).
Claim 3.1.G.
For each $i,j\in[t]$ with $i\neq j$, $\mathsf{d}\left(Z_{i},Q_{j}\right)\geq K$.
Proof.
Suppose that there exist $i,j\in[t]$ with $i\neq j$, and vertices $x\in V(Q^{\prime}_{j}),y\in Z_{i}$ such that $\mathsf{d}\left(x,y\right)<K$. By definition of $Q^{\prime}_{j}$, all its vertices are at distance at least $K$ from $Q_{i}$, so $y$ cannot be in $Q^{\prime}_{i}$, which includes $b_{i}$ and $b^{\prime}_{i}$ by ˜3.1.D. Hence, $y\in(V(P)\cap Z_{i})\setminus\{b_{i}\}$. Observe that, the $(y,b_{j})$-subpath of $P$ contains at least $5K+2$ antichain vertices of $A$, which is a contradiction due to ˜3.1.E. ∎
Now we prove Item˜6.
Claim 3.1.H.
@@
Suppose that there exists an integer $i\in[t]$, such that $\mathsf{d}\left(x,y\right)<K$ with $x\in T$ and $y\in Z_{i}$. The definition of $T$ implies that $x\in V(Q_{j})$ for some $j\in[t]$. By ˜3.1.G, it remains to consider the case when $i=j$. Recall that $c_{i}$ denotes the end-vertex of $Q^{\prime}_{i}$ that is different from $b_{i}$. From the definition of $T$ (and $T_{i}$), it follows that $\mathsf{d}\left(r,x\right)\leq\mathsf{d}\left(r,c_{i}\right)$.
Let $d^{\prime}_{i}$ denote the vertex of $Q_{i}$ which is adjacent to $c_{i}$ but does not lie in $Q^{\prime}_{i}$. The definition of $Q^{\prime}_{i}$ implies that there exists an integer $k\in[t]\setminus\{i\}$, such that $\mathsf{d}\left(d^{\prime}_{i},Q_{k}\right)<K$. Let $d^{\prime}_{k}$ denote a vertex of $Q_{k}$ with $\mathsf{d}\left(d^{\prime}_{i},d^{\prime}_{k}\right)<K$.
[…]
(a) (b)
Figure 7: Illustration of the proof of ˜3.1.H. Dashed lines indicate isometric paths of length at most $K$. (a) and (b) illustrate the cases when $\mathsf{d}\left(r,y\right)\leq\mathsf{d}\left(r,d^{\prime}_{i}\right)$, and when $\mathsf{d}\left(r,y\right)>\mathsf{d}\left(r,d^{\prime}_{i}\right)$, respectively. Dotted lines indicates that the two end-vertices of the line are at the same distance from $r$.
First, consider the case when $\mathsf{d}\left(r,y\right)\leq\mathsf{d}\left(r,d^{\prime}_{i}\right)$. See Figure˜7(a). Let $z$ be the vertex of $Q_{k}$ such that $\mathsf{d}\left(r,z\right)=\min\{\mathsf{d}\left(r,b^{\prime}_{k}\right),\mathsf{d}\left(r,d^{\prime}_{i}\right)\}$. We claim that, $\mathsf{d}\left(d^{\prime}_{k},z\right)\leq K$. When $z\neq b^{\prime}_{k}$, since $\mathsf{d}\left(d^{\prime}_{i},d^{\prime}_{k}\right)\leq K$, the vertex $z$ which is at same distance from $r$ as $d^{\prime}_{i}$ must lie at a distance at most $K$ from $d^{\prime}_{k}$. When $z=b_{k}$, $\mathsf{d}\left(r,b^{\prime}_{k}\right)\geq\mathsf{d}\left(r,c\right)$ and since $d^{\prime}_{k}$ lies in the $(z,r)$-isometric path, $\mathsf{d}\left(d^{\prime}_{k},z\right)\leq K$.
Let $y^{\prime}$ be the vertex of $Q_{i}$ such that $\mathsf{d}\left(r,y\right)=\mathsf{d}\left(r,y^{\prime}\right)$. Using similar arguments as above, it can be shown that $\mathsf{d}\left(y^{\prime},x\right)\leq K$. Now observe the following:
@@
$\leq 1+\mathsf{d}\left(b^{\prime}_{k},z\right)+\mathsf{d}\left(z,d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(d^{\prime}_{i},y^{\prime}\right)+\mathsf{d}\left(y^{\prime},x\right)+\mathsf{d}\left(x,y\right)$
$\leq 1+(\mathsf{d}\left(b^{\prime}_{k},z\right)+\mathsf{d}\left(d^{\prime}_{i},y^{\prime}\right))+\mathsf{d}\left(z,d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(y^{\prime},x\right)+\mathsf{d}\left(x,y\right)$
Observe that $\mathsf{d}\left(b^{\prime}_{k},z\right)=\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,z\right)$ and $\mathsf{d}\left(d^{\prime}_{i},y^{\prime}\right)=\mathsf{d}\left(r,d^{\prime}_{i}\right)-\mathsf{d}\left(r,y^{\prime}\right)$. Now consider two cases. If $z\neq b^{\prime}_{k}$, then $\mathsf{d}\left(r,z\right)=\mathsf{d}\left(r,d^{\prime}_{i}\right)$ and therefore,
$\mathsf{d}\left(b_{k},y\right)$ $\leq 1+(\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,z\right))+(\mathsf{d}\left(r,d^{\prime}_{i}\right)-\mathsf{d}\left(r,y^{\prime}\right))+\mathsf{d}\left(z,d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(y^{\prime},x\right)+\mathsf{d}\left(x,y\right)$
$\leq 1+(\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,d^{\prime}_{i}\right)+\mathsf{d}\left(r,d^{\prime}_{i}\right)-\mathsf{d}\left(r,y^{\prime}\right))+\mathsf{d}\left(z,d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(y^{\prime},x\right)+\mathsf{d}\left(x,y\right)$
$<1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,y\right)|+4K$